OpenStudy (anonymous):

if i were to provide all the possible rational zeros to f(x) = 9x^4 – 17x^3 + 2x^2 – 3x + 33 would i get : 1/9, 1/3, 11/3 ? when simplified

OpenStudy (anonymous):

make a fraction where numerator is all divisors of 33, denominator is all divisors of 9 there are a lot, and don't forget \[\pm\]

OpenStudy (anonymous):

\[\pm1,\pm3, \pm11, \pm 33,\pm \frac{1}{3},\pm\frac{11}{3},\pm\frac{1}{9},\pm\frac{11}{9}\] maybe even some i missed

OpenStudy (anonymous):

oh so i wouldnt simplify the fraction 33/9?

OpenStudy (anonymous):

yes